ganeshie8 (ganeshie8):

Find all integers \(n\) that satisfy the equation \[(n-1)^3 + n^3 + (n+1)^3 = (n+2)^3\] [Hint: Work with the equation obtained by replacing \(n\) by \(k+4\).]

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ (n-1)^3+n^3+(n+1)^3=(n+2)^3 \hspace{.33em}\\~\\ \implies \cdot\cdot\cdot\cdot \hspace{.33em}\\~\\ \implies \cdot\cdot\cdot\cdot \hspace{.33em}\\~\\ 2n^3-6n^2-6n-8=0 \hspace{.33em}\\~\\ n^3-3n^2-3n-4=0 \hspace{.33em}\\~\\ \color{red}{1}n^3-3n^2-3n-\color{blue}{4}=0 \hspace{.33em}\\~\\ \normalsize \text{by rational roots theorm} \hspace{.33em}\\~\\ \normalsize \text{possible roots are factors of 4 by factors of 1} \hspace{.33em}\\~\\ n=\{ \pm\dfrac{\color{blue}{4}}{\color{red}{1}},\pm\dfrac{\color{blue}{2}}{\color{red}{1}},\pm\dfrac{\color{blue}{1}}{\color{red}{1}}\}\hspace{.33em}\\~\\ \normalsize \text{by synthetic division} \hspace{.33em}\\~\\ 4|1 ~~-3~~-3~~-4\hspace{.33em}\\~\\ \underline{~~~- ~~~~4~~~~~~4~~~~~~~4}\hspace{.33em}\\~\\ ~~~~~1~~~~~1~~~~~~1~~~~~~0\hspace{.33em}\\~\\ \normalsize \text{n=4 satisfies it} \hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

Nice :) i didn't think of rational root thm before xD

OpenStudy (michele_laino):

Please note that, if I compute the third powers, i get this: \[n ^{3}-1-3n ^{2}+3n+n ^{3}+n ^{3}+1+3n ^{2}+3n=\] \[=n ^{3}+8+6n ^{2}+12n\] and after a simplification, I get: \[n ^{3}-3n ^{2}-3n-4=0\] which can be re-written as below: \[n ^{3}-1-3(n ^{2}+n+1)=0\] then we can write: \[[(n-1)-3](n ^{2}+n+1)-3(n ^{2}+n+1)=0\] now I factor out the quantity \[n ^{2}+n+1\] and I get: \[(n-4)(n ^{2}+n+1)=0\] which admits the only real solution n=4

ganeshie8 (ganeshie8):

Wow! that looks great xD

ganeshie8 (ganeshie8):

I'm thinking why the textbook gave that hint lol

ganeshie8 (ganeshie8):

both above methods look neat to me and they didnt use the hint, so that makes me feel the textbook hint is kinda unnecessary..

OpenStudy (mathmath333):

@Michele_Laino 's solution looks creative !

OpenStudy (michele_laino):

Thanks! @mathmath333

OpenStudy (michele_laino):

Thanks for your appreciation @ganeshie8

OpenStudy (anonymous):

hehehe well that hint was to reduce work :3 \((n+3)^3+(n+4)^3+(n+5)^3-(n+6)^3=0\\ \downarrow \\ (n^3+9n^2+27n+27)+(n^3+12n^2+48n+64)+\\(n^3+15n^2+75n+125)-(n^3+18n^2+108n+216)=0\\ \downarrow \\ 27+64+125-216=0\\ \downarrow \\ (n^3+9n^2+27n)+(n^3+12n^2+48n)+\\(n^3+15n^2+75n)-(n^3+18n^2+108n)=0\\ \downarrow \\ n[(n^2+9n+27)+(n^2+12n+48)+(n^2+15n+75)-(n^2+18n +108)]=0\\ n[(2n^2+18n+42)]=0\\ 2n(n^2+9n+21)\\ n=0 \text{ the real solution }\\ n^2+9n+21 \text{ has no real solution }\\\)

OpenStudy (anonymous):

and one thing i dint find this question in my book xD what is CHMP is it the appendix ?